# Active Learning with Variational Quantum Circuits for Quantum Process Tomography

### v1

@@ Introduction
In this section, we first analyze the inefficiencies in the crucial query step. Then, we propose a general framework for AL that utilizes the differences in informativeness among quantum states to optimize this step. Finally, we design various AL algorithms with the aim of offering choices suited to different scenarios.
Active Learning
@@
Considering that querying a quantum system is prohibitively costly since it requires multiple experiments, our goal is to learn the parameters $\boldsymbol{\theta}$ with error $\epsilon$ from training dataset $\{(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\}_{i=1}^{m}$, while minimizing $m$. However, in the query step, it is common to select a subset of quantum states randomly. This straightforward strategy typically requires a large number of labeled quantum states to train a VQC to achieve satisfactory performance. In fact, each quantum state contains different informativeness, with those having greater informativeness typically improving training more efficiently. This random selection strategy does not take this into account, resulting in training inefficiencies. To overcome this limitation, we focus on the informativeness of quantum states, aiming to select a subset of quantum states that improves the training most efficiently.
Acquisition Function.
@@
AL is a machine learning paradigm that promises to reduce annotation costs by intelligently selecting a subset of informative samples for labeling (Settles 2009). The selection of such a subset relies on a so-called acquisition function, which is key to AL. Specifically, the acquisition function evaluates the informativeness of each unlabeled sample and ranks them. The highest-ranking sample is then selected for labeling and subsequently added to the training dataset for updating the model. This process is repeated, gradually increasing the number of labeled samples.
Active Learning Framework.
Our framework for AL is illustrated in Fig. 2 and can be briefly described as follows:
  • Initialize: Initialize an unlabeled pool $D_{u}$ with the informationally complete set of quantum states, an empty labeled pool $D_{l}$, and a base model VQC.
@@
  • Evaluate: Use the acquisition function to evaluate and rank each quantum state in the unlabeled pool $D_{u}$.
  • Select: Select the highest-ranking quantum state $|\psi^{\star}\rangle$.
@@
  • Query: Query the unknown unitary process $\mathcal{U}$ to obtain the label for the selected quantum state.
  • Add: Add the newly labeled quantum state to the labeled pool $D_{l}$.
  • Update: Update the model with the newly incorporated labeled pool $D_{l}$.
@@
The process follows an iterative procedure in which quantum states are labeled sequentially to improve the model performance. At each AL step, the most informative quantum state is selected based on the acquisition function. This cycle is repeated until the labeling budget is exhausted or a satisfactory model performance is achieved.
@@
Based on various acquisition functions, current AL methods can be broadly classified into committee-based, uncertainty-based and diversity-based (Kumar and Gupta 2020). With the aim of offering choices suited to different scenarios, we design various AL algorithms to each category as follows.
Query-by-Committee
@@
Query-by-committee (QBC) is a widely used AL approach (Freund et al. 1997; Burbidge, Rowland, and King 2007). Its basic idea is to build a committee of models from existing labeled data, and then selects the sample for which the committee disagree most to label.
@@
We construct a committee of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, which are initialized with different parameters and trained on $D_{l}$. In each AL step, each member of the committee $\boldsymbol{\Theta}$ casts its vote based on their learned quantum process representations $\{\mathcal{C}(\boldsymbol{\theta}_{1}),\mathcal{C}(\boldsymbol{\theta}_{2}),\ldots,\mathcal{C}(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\}$. The committee select the quantum state for which all models disagree most on the predictions $\{|\psi(\boldsymbol{\theta}_{1})\rangle,|\psi(\boldsymbol{\theta}_{2})\rangle\},\ldots,|\psi(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\rangle\}$. A common strategy to calculate the level of disagreement is the variance. More precisely, for each unlabeled quantum state $|\psi\rangle\in D_{u}$, we compute the variance of the $n_{\mathrm{VQC}}$ individual predictions and then select the quantum state with the maximum variance, that is,
@@
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}d\left(|\psi(\boldsymbol{\theta}_{k})\rangle,|\psi(\boldsymbol{\Theta})\rangle\right),$ (4)
@@
where $|\psi({\boldsymbol{\Theta}})\rangle\propto\sum_{k=1}^{n_{\mathrm{VQC}}}|\psi({\boldsymbol{\theta}_{k}})\rangle$, and $d$ represents the Euclidean distance. The pseudo-code for QBC is shown in Algorithm 1.
@@
1:  Construct a committee of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
@@
4:        $|\psi(\boldsymbol{\theta}_{k})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta}_{k})|\psi\rangle$
@@
6:     Calculate the variance of $\{|\psi(\boldsymbol{\theta}_{k})\rangle\}_{k=1}^{n_{\mathrm{VQC}}}$ using Eq. (4)
@@
Expected Model Change Maximization
Expected model change maximization (EMCM) is also a popular AL approach that aims to select the sample resulting in the largest change to the current model (Cai, Zhang, and Zhou 2013). The model change is measured as the difference between the current model parameters and the updated parameters after training with the enlarged training dataset. Inspired by Stochastic Gradient Descent (SGD) update rule, the change is estimated as the gradient of the loss with respect to an unlabeled sample.
Specifically, we employ SGD to update the parameters $\boldsymbol{\theta}$. The parameters $\boldsymbol{\theta}$ are updated iteratively according to the negative gradient descent of the loss $\ell(\boldsymbol{\theta})$ with respect to each training data $(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\in D_{l}$:
$\boldsymbol{\theta}:=\boldsymbol{\theta}-\alpha\nabla_{\boldsymbol{\theta}}\ell_{|\psi_{i}\rangle}(\boldsymbol{\theta})\quad\text{for }i=1,2,\ldots,m,$ (5)
@@
$\mathcal{L}_{D_{l}^{+}}=\frac{1}{|D_{l}^{+}|}\left[\sum_{|\psi_{i}\rangle\in D_{l}}\left\||\psi_{i}\rangle-|\psi_{i}^{\mathrm{ideal}}\rangle\right\|^{2}+\underbrace{\left\||\psi\rangle-|\psi^{\mathrm{ideal}}\rangle\right\|^{2}}_{=:\ell_{|\psi\rangle}(\boldsymbol{\theta})}\right].$ (6)
Thus, the model obtained by minimizing the loss function is changed due to the inclusion of the new training data $(|\psi\rangle,|\psi^{\mathrm{ideal}}\rangle)$. We measure the model change as the parameter change using the gradient of the loss at $|\psi\rangle$. The derivative of the loss $\ell_{|\psi\rangle}(\boldsymbol{\theta})$ with respect to the parameters $\theta_{i}$ at $|\psi\rangle$ is formulated as:
$\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\theta_{i}}$ $=\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}$
$=\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\left[\frac{1}{2}\left(\mathcal{O}(\boldsymbol{\theta}_{i}^{+})-\mathcal{O}(\boldsymbol{\theta}_{i}^{-})\right)\right],$ (7)
where $\mathcal{O}(\boldsymbol{\theta})=\mathrm{Re}\left(\left\langle\psi^{\mathrm{ideal}}\mid\psi(\boldsymbol{\theta})\right\rangle\right)$. Since the true label $|\psi^{\mathrm{ideal}}\rangle$ is actually unknown before querying in practice, we construct an ensemble $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$ to estimate the prediction $\{|\psi(\boldsymbol{\theta}_{1})\rangle,|\psi(\boldsymbol{\theta}_{2})\rangle,\ldots,|\psi(\boldsymbol{\theta}_{n_{\mathrm{VQC}}})\rangle\}$, and use the expected model change to approximate the true model change, that is:
$\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\theta_{i}}$ $=\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}\frac{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}{\partial\theta_{i}}$
$=\frac{1}{2n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\boldsymbol{\theta})}{\partial\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})}\left[\hat{\mathcal{O}}_{k}(\boldsymbol{\theta}_{i}^{+})-\hat{\mathcal{O}}_{k}(\boldsymbol{\theta}_{i}^{-})\right],$ (8)
where $\hat{\mathcal{O}}_{k}(\boldsymbol{\theta})=\mathrm{Re}\left(\left\langle\psi(\boldsymbol{\theta}_{k})\mid\psi(\boldsymbol{\theta})\right\rangle\right)$ represents an estimate of $\mathcal{O}$.
@@
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\left\|\nabla_{\boldsymbol{\theta}}\ell_{|\psi\rangle}(\boldsymbol{\theta})\right\|.$ (9)
@@
Input: Labeled quantum states set $D_{l}$, unlabeled quantum states set $D_{u}$, base model $\mathcal{C}(\boldsymbol{\theta})$ trained on $D_{l}$, number of ensemble models $n_{\mathrm{VQC}}$
@@
1:  Construct an ensemble of $n_{\mathrm{VQC}}$ models $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
@@
3:     $|\psi(\boldsymbol{\theta})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta})|\psi\rangle$
@@
5:        $|\psi(\boldsymbol{\theta}_{k})\rangle\leftarrow\mathcal{C}(\boldsymbol{\theta}_{k})|\psi\rangle$
@@
7:     Calculate the gradient norm $\left\|\nabla_{\boldsymbol{\theta}}\ell_{|\psi\rangle}(\boldsymbol{\theta})\right\|$ using Eq.(8)
@@
Greedy Sampling
Instead of finding the most informative sample based on the learned model, as in QBC and EMCM, greedy sampling (GS) is a model-free approach that selects the sample based on its geometric characteristics in the feature space (Yu and Kim 2010). Its basic idea is to greedily select an unlabeled sample that is located far from the previously labeled samples, thereby ensuring that the selected samples are scattered across the entire input space rather than concentrating in a small local region. Clearly, this approach allows for selecting a set of samples at the beginning of the AL process, making it easy to implement by avoiding the need to evaluate each unlabeled sample and update the model at each AL step.

### v2

@@ III Method
In this section, we first analyze the inefficiencies in the crucial query step. To address this issue, we then propose a general AL framework that explicitly utilizes the differences in informativeness among quantum states to optimize the query step. Finally, we design various AL algorithms with the aim of offering choices suited to different scenarios. Notably, we emphasize our AL framework is general and can in principle be applied in the query step of any QPT method, not just limited to the VQCs considered here.
III.1 Active Learning
@@
Considering that querying a quantum system is prohibitively costly since it requires multiple experiments, our goal is to learn the parameters $\bm{\theta}$ with error $\epsilon$ from training dataset $\{(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\}_{i=1}^{m}$, while minimizing the number of labeled quantum states $m$ (i.e., queries to the unknown process $\mathcal{U}$). However, in the query step, it is common to select a subset of quantum states randomly. This straightforward strategy generally requires a large number of labeled quantum states to train a VQC to achieve satisfactory performance. As the system scale increases, this challenge becomes more pronounced—it is always the case that there are insufficient labeled quantum states to train a VQC with desirable performance for high-quality reconstruction. In fact, each quantum state contains different informativeness, with those having greater informativeness typically improving training more efficiently. This random selection strategy does not take this into account, resulting in training inefficiencies. To overcome this limitation, we focus on the informativeness of quantum states, aiming to intelligently select a subset of quantum states that improves the training most efficiently.
@@
Acquisition Function.— AL is a machine learning paradigm that promises to reduce annotation costs by intelligently selecting a subset of informative samples for labeling [17]. The selection of such a subset relies on a so-called acquisition function. This function plays a central role in AL. It operates by evaluating samples based on criteria such as committee disagreement, prediction uncertainty, or diversity, which essentially estimate the informativeness of samples from different perspectives [24]. Specifically, the acquisition function evaluates the informativeness of each unlabeled sample and ranks them accordingly. The highest-ranking sample is then selected for labeling and subsequently added to the training dataset for updating the model. This process is repeated, gradually increasing the number of labeled samples and improving the model’s performance.
Active Learning Framework.— A schematic overview of our AL framework is provided in Fig. 1. It consists of the following steps:
  1. 1.
    Initialize: Begin by preparing an unlabeled dataset $D_{u}$ that consists of an informationally complete set of quantum states (i.e., a set of quantum states sufficient to fully characterize the quantum process). Also, initialize an empty labeled pool $D_{l}$. It will be gradually expanded as labeled quantum states are added during the AL process. Finally, set up a base VQC model $\mathcal{C}(\bm{\theta})$, which will be iteratively trained to learn the unknown quantum process $\mathcal{U}$.
  2. 2.
@@
    Evaluate: Use the acquisition function to evaluate the informativeness of each quantum state in the unlabeled pool $D_{u}$. Then, rank all the states according to these information values, thereby effectively identifying the most informative quantum state for labeling.
  3. 3.
    Select: Select the highest-ranking quantum state $|\psi^{\star}\rangle$ from the unlabeled pool $D_{u}$. This carefully chosen quantum state is expected to contribute most efficiently to improving the VQC model’s performance.
  4. 4.
@@
    Query: Query the unknown unitary process $\mathcal{U}$ by inputting the selected quantum state $|\psi^{\star}\rangle$, and obtain the corresponding output $\mathcal{U}|\psi^{\star}\rangle$, which serves as its true label.
  5. 5.
    Add: Add the newly labeled quantum state $(|\psi^{\star}\rangle,\mathcal{U}|\psi^{\star}\rangle)$ to the labeled pool $D_{l}$. This step enables the labeled pool $D_{l}$ to gradually expand throughout the AL process, thereby improving the training of the VQC model $\mathcal{C}(\bm{\theta})$ to better approximate the unknown quantum process $\mathcal{U}$.
  6. 6.
    Update: Retrain the VQC model $\mathcal{C}(\bm{\theta})$ using the updated labeled pool $D_{l}$, which now includes the newly added quantum state.
@@
The process follows an iterative procedure in which quantum states are sequentially labeled to improve model performance. At each AL step, the most informative quantum state is selected based on the acquisition function. This AL cycle of EvaluateSelectQueryAddUpdate repeats until the labeling budget is exhausted or a satisfactory model performance is achieved.
@@
Based on various acquisition functions, current AL methods can be broadly classified into committee-based, uncertainty-based and diversity-based [24]. As mentioned earlier, these approaches essentially estimate the informativeness of samples from different perspectives. With the aim of offering choices suited to different scenarios, we design various AL algorithms to each category as follows.
III.2 Query-by-Committee
@@
Query-by-committee (QBC) is a widely used AL approach [27, 28]. Its basic idea is to build a committee of models from existing labeled data, and then selects the sample for which the committee disagree most to label. This is because the degree of disagreement among the committee members can serve as an estimate of the information value.
@@
We construct a committee of $n_{\mathrm{VQC}}$ models $\bm{\Theta}=\{\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{n_{\mathrm{VQC}}}\}$, which are initialized with different parameters and trained on $D_{l}$. In each AL step, each member of the committee $\bm{\Theta}$ casts its vote based on their learned quantum process representations $\{\mathcal{C}(\bm{\theta}_{1}),\mathcal{C}(\bm{\theta}_{2}),\ldots,\mathcal{C}(\bm{\theta}_{n_{\mathrm{VQC}}})\}$. The committee select the quantum state for which all models disagree most on the predictions $\{|\psi(\bm{\theta}_{1})\rangle,|\psi(\bm{\theta}_{2})\rangle\},\ldots,|\psi(\bm{\theta}_{n_{\mathrm{VQC}}})\rangle\}$. A common strategy to calculate the level of disagreement is the variance. More precisely, for each unlabeled quantum state $|\psi\rangle\in D_{u}$, we compute the variance of the $n_{\mathrm{VQC}}$ individual predictions and then select the quantum state with the maximum variance, that is,
@@
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}d\left(|\psi(\bm{\theta}_{k})\rangle,|\psi(\bm{\Theta})\rangle\right),$ (4)
@@
where $|\psi({\bm{\Theta}})\rangle\propto\sum_{k=1}^{n_{\mathrm{VQC}}}|\psi({\bm{\theta}_{k}})\rangle$, and $d$ represents the Euclidean distance. The pseudo-code for QBC is shown in Algorithm 1.
@@
1:  Construct a committee of $n_{\mathrm{VQC}}$ models $\bm{\Theta}=\{\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
@@
4:        $|\psi(\bm{\theta}_{k})\rangle\leftarrow\mathcal{C}(\bm{\theta}_{k})|\psi\rangle$
@@
6:     Calculate the variance of $\{|\psi(\bm{\theta}_{k})\rangle\}_{k=1}^{n_{\mathrm{VQC}}}$ using Eq. (4)
@@
III.3 Expected Model Change Maximization
Expected model change maximization (EMCM) is also a popular AL approach that aims to select the sample resulting in the largest change to the current model [29]. The underlying rationale is that samples causing significant changes to the model are highly informative. They help reveal regions of the data distribution that the current model has not yet fully captured. For simplicity, the model change is typically measured as the difference between the current model parameters and the updated parameters after training with the enlarged training dataset. Inspired by Stochastic Gradient Descent (SGD) update rule, the change is estimated as the gradient of the loss with respect to an unlabeled sample.
Specifically, we employ SGD to update the parameters $\bm{\theta}$. The parameters $\bm{\theta}$ are updated iteratively according to the negative gradient descent of the loss $\ell(\bm{\theta})$ with respect to each training data $(|\psi_{i}\rangle,|\psi_{i}^{\mathrm{ideal}}\rangle)\in D_{l}$:
$\bm{\theta}:=\bm{\theta}-\alpha\nabla_{\bm{\theta}}\ell_{|\psi_{i}\rangle}(\bm{\theta})\quad\text{for }i=1,2,\ldots,m,$ (5)
@@ III.3 Expected Model Change Maximization
$\mathcal{L}_{D_{l}^{+}}=\frac{1}{|D_{l}^{+}|}\left[\sum_{|\psi_{i}\rangle\in D_{l}}\left\||\psi_{i}\rangle-|\psi_{i}^{\mathrm{ideal}}\rangle\right\|^{2}+\underbrace{\left\||\psi\rangle-|\psi^{\mathrm{ideal}}\rangle\right\|^{2}}_{=:\ell_{|\psi\rangle}(\bm{\theta})}\right].$ (6)
Thus, the model obtained by minimizing the loss function is changed due to the inclusion of the new training data $(|\psi\rangle,|\psi^{\mathrm{ideal}}\rangle)$. We measure the model change as the parameter change using the gradient of the loss at $|\psi\rangle$. The derivative of the loss $\ell_{|\psi\rangle}(\bm{\theta})$ with respect to the parameters $\theta_{i}$ at $|\psi\rangle$ is formulated as:
$\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\theta_{i}}$ $=\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\mathcal{O}(\bm{\theta})}\frac{\partial\mathcal{O}(\bm{\theta})}{\partial\theta_{i}}$ (7)
$=\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\mathcal{O}(\bm{\theta})}\left[\frac{1}{2}\left(\mathcal{O}(\bm{\theta}_{i}^{+})-\mathcal{O}(\bm{\theta}_{i}^{-})\right)\right],$
where $\mathcal{O}(\bm{\theta})=\mathrm{Re}\left(\left\langle\psi^{\mathrm{ideal}}\mid\psi(\bm{\theta})\right\rangle\right)$. Since the true label $|\psi^{\mathrm{ideal}}\rangle$ is actually unknown before querying in practice, we construct an ensemble $\bm{\Theta}=\{\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{n_{\mathrm{VQC}}}\}$ to estimate the prediction $\{|\psi(\bm{\theta}_{1})\rangle,|\psi(\bm{\theta}_{2})\rangle,\ldots,|\psi(\bm{\theta}_{n_{\mathrm{VQC}}})\rangle\}$, and use the expected model change to approximate the true model change, that is:
$\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\theta_{i}}$ $=\frac{1}{n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\hat{\mathcal{O}}_{k}(\bm{\theta})}\frac{\partial\hat{\mathcal{O}}_{k}(\bm{\theta})}{\partial\theta_{i}}$
$=\frac{1}{2n_{\mathrm{VQC}}}\sum_{k=1}^{n_{\mathrm{VQC}}}\frac{\partial\ell_{|\psi\rangle}(\bm{\theta})}{\partial\hat{\mathcal{O}}_{k}(\bm{\theta})}\left[\hat{\mathcal{O}}_{k}(\bm{\theta}_{i}^{+})-\hat{\mathcal{O}}_{k}(\bm{\theta}_{i}^{-})\right],$ (8)
where $\hat{\mathcal{O}}_{k}(\bm{\theta})=\mathrm{Re}\left(\left\langle\psi(\bm{\theta}_{k})\mid\psi(\bm{\theta})\right\rangle\right)$ represents an estimate of $\mathcal{O}$.
@@
$|\psi^{\star}\rangle=\underset{|\psi\rangle\in D_{u}}{\operatorname{argmax}}\left\|\nabla_{\bm{\theta}}\ell_{|\psi\rangle}(\bm{\theta})\right\|.$ (9)
@@
Input: Labeled quantum states set $D_{l}$, unlabeled quantum states set $D_{u}$, base model $\mathcal{C}(\bm{\theta})$ trained on $D_{l}$, number of ensemble models $n_{\mathrm{VQC}}$
@@
1:  Construct an ensemble of $n_{\mathrm{VQC}}$ models $\bm{\Theta}=\{\bm{\theta}_{1},\bm{\theta}_{2},\ldots,\bm{\theta}_{n_{\mathrm{VQC}}}\}$, trained on $D_{l}$
@@
3:     $|\psi(\bm{\theta})\rangle\leftarrow\mathcal{C}(\bm{\theta})|\psi\rangle$
@@
5:        $|\psi(\bm{\theta}_{k})\rangle\leftarrow\mathcal{C}(\bm{\theta}_{k})|\psi\rangle$
@@
7:     Calculate the gradient norm $\left\|\nabla_{\bm{\theta}}\ell_{|\psi\rangle}(\bm{\theta})\right\|$ using Eq.(III.3)
@@
III.4 Greedy Sampling
Instead of finding the most informative sample based on the learned model, as in QBC and EMCM, greedy sampling (GS) is a model-free approach that selects the sample based on its geometric characteristics in the feature space [30]. Its basic idea is to greedily select an unlabeled sample that is located far from the previously labeled samples, thereby ensuring that the selected samples are scattered across the entire input space rather than concentrating in a small local region. This strategy effectively maximizes coverage and diversity of the selected samples, which are more likely to capture varied and complementary information. Clearly, GS allows for selecting a set of samples at the beginning of the AL process, making it easy to implement by avoiding the need to evaluate each unlabeled sample and update the model at each AL step.